An airplane flies from New Orleans to Atlanta at an average rate of 320 mph. The airplane then returns at an average rate of 280 mph. The total travel time is 3 hours. How long did it take the plane to make the return trip.
Answer by scott8148(6628)

distance = rate X time...the distances are the same so (rate out)X(time out)=(rate back)X(time back)...let x=(time back)

(time out)+(time back)=3...so (time out)=3-(time back)

320*(3-x)=280*x...or 960-320x=280x...or 960=600x...so x=1.6 hrs...or 1 hr 36 min